# Beyond Discretization: Learning the Optimal Solution Path

### v2

@@ 6 Numerical Experiments
We next empirically compare i) our approach using a fixed, large basis (denoted LSP for “Learning the Solution Path”) ii) our adaptive Algorithm 1 (denoted ALSP for “Adaptive LSP” and iii) uniform discretization, a natural benchmark. We aim to show that both our approaches not only outperform the benchmark but that the qualitative insights from our theoretical results hold for more general optimization procedures than constant step-size SGD. We also provide preliminary results on how the choice of basis affects performance.
We choose uniform discretization for our benchmark in lieu of other schemes (like geometric spacing) because i) it matches the theoretical lower bound from (Ndiaye et al., 2019) and ii) we see it as most intuitive for learning the solution path with small uniform error. Specifically, for various $\epsilon$, we consider a uniform spacing of size $\sqrt{\epsilon}$ in each dimension of $\lambda$, and run (warm-started) gradient descent for $O(\log(1/\epsilon))$ iterations. The constant here is calibrated in an oracle fashion to achieve a solution-path error of $O(\epsilon)$ (see Section B.1 for details) giving the benchmark a small advantage.
@@ 6.1 Weighted Binary Classification
$h(\boldsymbol{\theta},\lambda)=(1-\lambda)l_{\text{pos}}(\boldsymbol{\theta})+\lambda l_{\text{neg}}(\boldsymbol{\theta})+0.125\|\boldsymbol{\theta}\|^{2},$
where $l_{\text{pos}}(\boldsymbol{\theta})$ and $l_{\text{neg}}(\boldsymbol{\theta})$ denote the log-likelihood on the positive and negative classes respectively. Specifically, letting $(\boldsymbol{x}_{i},y_{i})\in\mathbb{R}^{45}\times\{0,1\}$ for $i=1,\ldots,n$ denote the data,
$l_{\text{pos}}(\boldsymbol{\theta})=\frac{1}{|\{i:\ y_{i}=1\}|}\sum_{i:y_{i}=1}\log(1+e^{(-2y_{i}+1)\boldsymbol{x}_{i}^{\top}\boldsymbol{\theta}}),$
and similarly for $l_{\text{neg}}(\boldsymbol{\theta})$.
@@
The ground truth $\boldsymbol{\theta}^{*}(\lambda)$ is computed via $5000$ iterations of (warm-started) gradient descent over a uniform grid of $2^{10}$ points. Solution path error is approximated by the uniform error over this grid.
For LSP, we run SGD using torch.optim.SGD. In lieu of a constant learning rate, we dynamically reduce the learning rate according to the Distance Diagnostic of Pesme et al. (2020, Section 4). This dynamic updating is more reflective of practice. We use the suggested parameters from Pesme et al. (2020) with the exception of $q$, which we tune by examining the performance after $200$ iterations. (We take $q=1.3$ for both the Legendre and Jacobi bases.) For ALSP, we initialize the algorithm with $5$ polynomials and stop it after reaching $12$ polynomials ($p=5$ to $12$).
Figure 2: LSP and ALSP for Weighted Binary Classification. (a) Compares methods using Jacobi polynomials. (b) Error in solution path as a function of $\lambda$ using Jacobi polynomials. (c) and (d) compare $\boldsymbol{\theta}_{3}(\lambda)$ across $\lambda$ in the Legendre and Jacobi bases respectively with $7$ polynomials.
@@
Figure 2 panel (b) shows that the largest errors occur almost periodically at a few places across $\Lambda=[0,1]$, matching the prediction by the equioscillation theorem.
@@ 6.2 Portfolio Allocation
We next consider a portfolio allocation problem calibrated to real data where $\boldsymbol{\theta}\in\mathbb{R}^{d}$ represents the weights on $d=10$ different asset classes. Namely, let $\mu\in\mathbb{R}^{d}$ and $\Sigma\in\mathbb{R}^{d\times d}$ be the mean and covariance matrix of the returns of the different asset classes. We fit these parameters to the monthly return data from Aug. 2014 to July 2024 using the Fama-French 10 Industry index dataset. [^1] https://mba.tuck.dartmouth.edu/pages/faculty/ken.french/data_library.html
@@
$\min_{\boldsymbol{\theta}}$ $\quad-\lambda_{1}\mu^{\top}\boldsymbol{\theta}+\lambda_{2}\boldsymbol{\theta}^{\top}\Sigma\boldsymbol{\theta}+\bar{\ell}_{1}(\boldsymbol{\theta}).$
@@
Ground truth $\boldsymbol{\theta}^{*}(\lambda)$ is computed over a $100\times 100$ grid.
We focus on bivariate-Legendre polynomials for our basis, scaled and shifted to be orthogonal to uniform distribution on $[0,1]\times[.2,1]$. We initialize with ALSP with 2 polynomials in each dimension, and iterate by adding one polynomial in each dimension, so that $p=4,9,16,25$, after which we stop.
@@
Figure 3: LSP and ALSP for Portfolio Allocation. Compares methods using the first $p=4,9,16$ bivariate-Legendre polynomials.
As in our previous experiment, we see that LSP rapidly converges to an irreducible error and then plateaus. By contrast, ALSP seems to make continued progress as we add polynomials. Both method substantively outperform uniform discretization.
The improved performance over uniform discretization is partially attributable to the increased dimension of $\Lambda$ (because discretization suffers the curse of dimensionality), but is also because (traditionally) discretization interpolates solutions in a piecewise constant fashion. In this example, $\boldsymbol{\theta}^{*}(\lambda)$ is very smooth. See Figure 5 in Appendix. Hence, polynomial can learn to interpolate values very fast, while uniform discretization needs a great deal more resolution.
@@ 7 Conclusion
We propose a new method for learning the optimal solution path of a family of problems by reframing it as a single stochastic optimization problem over a linear combination of pre-specified basis functions. Compared to discretization schemes, our method offers flexibility and scalability by taking a global perspective on the solution path. When using sufficiently large bases, we prove our problem satisfies a certain relaxed weak-growth condition that strongly allows us to solve the single optimization problm very efficiently. Theoretical results in special cases and numerical experiments in more general settings support these findings.
Future research might more carefully examine the interplay between the parameterization of the family of problems and choice of basis. One might also consider other universal function approximators (like deep neural networks or forests of trees) within this context.
@@ A.1 Proofs from Section 2
Suppose by contradiction that $c=0$. Then there exists a corresponding eigenvector $\boldsymbol{v}\in\mathbb{R}^{p}$ with $\boldsymbol{v}\neq\boldsymbol{0}$ such that
$\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\boldsymbol{\Phi}(\tilde{\lambda})^{\top}\boldsymbol{\Phi}(\tilde{\lambda})\right]\boldsymbol{v}=\boldsymbol{0}\implies\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\|\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{v}\|^{2}_{2}\right]=0.$
This further implies that $\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{v}=0$ almost everywhere. By the linear independence assumption in Assumption 2.2, we then must have $\boldsymbol{v}=\boldsymbol{0}$, but this is a contradiction since $\boldsymbol{v}$ is an eigenvector.
@@
$c\leq\sigma_{\max}\left\{\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\boldsymbol{\Phi}(\tilde{\lambda})^{\top}\boldsymbol{\Phi}(\tilde{\lambda})\right]\right\}\leq\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\sigma_{\max}\left\{\boldsymbol{\Phi}(\tilde{\lambda})^{\top}\boldsymbol{\Phi}(\tilde{\lambda})\right\}\right]\leq C,$
@@
$f(\boldsymbol{\beta},\lambda)~{}:=~{}h(\boldsymbol{\Phi}(\lambda)\boldsymbol{\beta},\lambda),\ \ \ \ F(\boldsymbol{\beta})~{}:=~{}\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[f(\boldsymbol{\beta},\tilde{\lambda})\right]=\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta},\tilde{\lambda})\right].$
Before proceeding, we establish smoothness and strong convexity of these functions.
@@
Fix an arbitrary $\lambda\in\Lambda$. Then, for any $\boldsymbol{\beta}$, $\bar{\boldsymbol{\beta}}\in\mathbb{R}^{p}$, smoothness of $f(\cdot,\lambda)$ is certified by
$\|\nabla f(\boldsymbol{\beta},\lambda)-\nabla f(\bar{\boldsymbol{\beta}},\lambda)\|$ $=\|\nabla_{\boldsymbol{\beta}}h(\boldsymbol{\Phi}(\lambda)\boldsymbol{\beta},\lambda)-\nabla_{\boldsymbol{\beta}}h(\boldsymbol{\Phi}(\lambda)\bar{\boldsymbol{\beta}},\lambda)\|$
@@
$\ \leq CL\|\boldsymbol{\beta}-\bar{\boldsymbol{\beta}}\|$
for any $\boldsymbol{\beta}$, $\bar{\boldsymbol{\beta}}\in\mathbb{R}^{d}$. To obtain the same conclusions for $F(\cdot)$, we may take expectation over $\lambda\sim\mathbb{P}_{\lambda}$ on the above inequalities, and invoke the linearity of expectations as well as the property that
@@
Next, we verify the strong convexity of $F(\cdot)$. Fixing $\boldsymbol{\beta}$, $\bar{\boldsymbol{\beta}}\in\mathbb{R}^{d}$, for any $\lambda\in\Lambda$, define $\boldsymbol{\theta}(\lambda):=\boldsymbol{\Phi}(\lambda)\boldsymbol{\beta}$, $\bar{\boldsymbol{\theta}}(\lambda):=\boldsymbol{\Phi}(\lambda)\bar{\boldsymbol{\beta}}$. Then, by strong convexity of $h$, we have
$h(\bar{\boldsymbol{\theta}}(\lambda),\lambda)\geq h(\boldsymbol{\theta}(\lambda),\lambda)+\nabla_{\boldsymbol{\theta}}h(\boldsymbol{\theta}(\lambda),\lambda)^{\top}\left(\bar{\boldsymbol{\theta}}(\lambda)-\boldsymbol{\theta}(\lambda)\right)+\frac{\mu}{2}\|\bar{\boldsymbol{\theta}}(\lambda)-\boldsymbol{\theta}(\lambda)\|^{2}.$
@@
$F(\bar{\boldsymbol{\beta}})$ $=\mathbb{E}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\bar{\boldsymbol{\beta}},\tilde{\lambda})\right]$
$\ \geq\mathbb{E}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta},\tilde{\lambda})\right]+\mathbb{E}\left[\nabla_{\boldsymbol{\theta}}h(\boldsymbol{\theta}(\tilde{\lambda}),\tilde{\lambda})^{\top}\left(\boldsymbol{\Phi}(\tilde{\lambda})(\bar{\boldsymbol{\beta}}-\boldsymbol{\beta})\right)\right]+\frac{\mu}{2}\mathbb{E}\left[\|\boldsymbol{\Phi}(\tilde{\lambda})(\bar{\boldsymbol{\beta}}-\boldsymbol{\beta})\|^{2}\right]$
$\ =F(\boldsymbol{\beta})+\mathbb{E}\left[\boldsymbol{\Phi}(\tilde{\lambda})^{\top}\nabla_{\boldsymbol{\theta}}h(\boldsymbol{\theta}(\tilde{\lambda}),\tilde{\lambda})\right]^{\top}(\bar{\boldsymbol{\beta}}-\boldsymbol{\beta})+\frac{\mu}{2}(\bar{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\top}\mathbb{E}\left[\boldsymbol{\Phi}(\tilde{\lambda})^{\top}\boldsymbol{\Phi}(\tilde{\lambda})\right](\bar{\boldsymbol{\beta}}-\boldsymbol{\beta})$
$\ \geq F(\boldsymbol{\beta})+\nabla F(\boldsymbol{\beta})^{\top}(\bar{\boldsymbol{\beta}}-\boldsymbol{\beta})+\frac{c\mu}{2}(\bar{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\top}(\bar{\boldsymbol{\beta}}-\boldsymbol{\beta})$
@@
First, we prove that $\epsilon_{\rm sp}^{\ast}$ is attained at some $\boldsymbol{\beta}_{\rm sp}^{\ast}\in\mathbb{R}^{p}$. Observe that $\epsilon_{\rm sp}(\cdot)$ is continuous. Thus, its level set $M(1):=\{\boldsymbol{\beta}\in\mathbb{R}^{p}:\epsilon_{\rm sp}(\boldsymbol{\beta})\leq 1\}$ is closed. Furthermore, $M(1)$ is bounded. To see this, for any $\boldsymbol{\beta}\in M(1)$, by strong convexity of $F$,
$\frac{c\mu}{2}\|\boldsymbol{\beta}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}$ $~{}\leq~{}F(\boldsymbol{\beta})-F(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast})$
$~{}=~{}\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta},\tilde{\lambda})-h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta}_{\mathrm{avg}}^{\ast},\tilde{\lambda})\right]$
$~{}\leq~{}\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta},\tilde{\lambda})-h(\boldsymbol{\theta}^{*}(\tilde{\lambda}),\tilde{\lambda})\right]$
$~{}\leq~{}\epsilon_{\rm sp}(\boldsymbol{\beta})$ (4)
@@
Since $\boldsymbol{\beta}$ is arbitrary, for any $\boldsymbol{\beta}$, $\bar{\boldsymbol{\beta}}\in M(1)$,
$\|\boldsymbol{\beta}-\bar{\boldsymbol{\beta}}\|~{}\leq~{}\|\boldsymbol{\beta}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|+\|\bar{\boldsymbol{\beta}}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|~{}\leq~{}\frac{2\sqrt{2}}{\sqrt{c\mu}}.$
Applying the Weierstrass Theorem, there exists $\boldsymbol{\beta}_{\rm sp}^{\ast}\in M(1)\subset\mathbb{R}^{p}$ s.t.
$\epsilon_{\rm sp}(\boldsymbol{\beta}_{\rm sp}^{\ast})~{}=~{}\epsilon_{\rm sp}^{\ast}.$
@@
$h(\boldsymbol{\Phi}(\lambda)\boldsymbol{\beta},\lambda)-h(\boldsymbol{\theta}^{*}(\lambda),\lambda)\ \leq\ \frac{L}{2}\|\boldsymbol{\Phi}(\lambda)\boldsymbol{\beta}-\boldsymbol{\theta}^{*}(\lambda)\|^{2}$ (5)

### v3

@@ 6 NUMERICAL EXPERIMENTS
We next empirically compare i) our approach using a fixed, large basis (denoted LSP for “Learning the Solution Path”) ii) our adaptive Algorithm 1 (denoted ALSP for “Adaptive LSP” and iii) uniform discretization, a natural benchmark. We aim to show that both our approaches not only outperform the benchmark but that the qualitative insights from our theoretical results hold for more general optimization procedures than constant step-size SGD. We also provide preliminary results on how the choice of basis affects performance. Our repository can be found at https://github.com/Cumberkid/Learning-the-Optimal-Solution-Path.
We choose uniform discretization to be our benchmark in lieu of other schemes (like geometric spacing) because i) it matches the theoretical lower bound from (Ndiaye et al., 2019) and ii) we see it as most intuitive for learning the solution path with small uniform error. Specifically, for various $\epsilon$, we consider a uniform spacing of size $\sqrt{\epsilon}$ in each dimension of $\lambda$, and run (warm-started) gradient descent for $O(\log(1/\epsilon))$ iterations. The constant hidden by big “Oh” here is calibrated in an oracle fashion to achieve a solution-path error of $O(\epsilon)$ (see Section B.1 for details) giving the benchmark a small advantage.
@@ 6.1 Weighted Binary Classification
$h(\bm{\theta},\lambda)=(1-\lambda)l_{\text{pos}}(\bm{\theta})+\lambda l_{\text{neg}}(\bm{\theta})+0.125\|\bm{\theta}\|^{2},$
where $l_{\text{pos}}(\bm{\theta})$ and $l_{\text{neg}}(\bm{\theta})$ denote the negative log-likelihood on the positive and negative classes respectively. Specifically, letting $(\bm{x}_{i},y_{i})\in\mathbb{R}^{45}\times\{0,1\}$ for $i=1,\ldots,n$ denote the data,
$l_{\text{pos}}(\bm{\theta})=\frac{1}{|\{i:\ y_{i}=1\}|}\sum_{i:y_{i}=1}\log(1+e^{(-2y_{i}+1)\bm{x}_{i}^{\top}\bm{\theta}}),$
and similarly for $l_{\text{neg}}(\bm{\theta})$.
@@
The ground truth $\bm{\theta}^{*}(\lambda)$ is computed via $5000$ iterations of (warm-started) gradient descent over a uniform grid of $2^{10}$ points. Solution path error is approximated by the uniform error over this grid.
For LSP, we run SGD using torch.optim.SGD. In lieu of a constant learning rate, we dynamically reduce the learning rate according to the Distance Diagnostic of Pesme et al. (2020, Section 4). This dynamic updating is more reflective of practice. We use the suggested parameters from Pesme et al. (2020) with the exception of $\mathtt{q}$ [^1] Note that $\mathtt{q}$, defined as a diagnostic frequency control parameter in Pesme et al. (2020, Section 4), is different from $q$, the number of polynomials in our paper., which we tune by examining the performance after $200$ iterations. (We take $\mathtt{q}=1.3$ for both the Legendre and Jacobi bases.) For ALSP, we initialize the algorithm with $5$ polynomials and stop it after reaching $12$ polynomials ($q=5$ to $12$).
Figure 2: LSP and ALSP for Weighted Binary Classification. (a) Compares methods using Jacobi polynomials. (b) Error in solution path as a function of $\lambda$ using Jacobi polynomials. (c) and (d) compare $\bm{\theta}_{3}(\lambda)$ across $\lambda$ in the Legendre and Jacobi bases respectively with $7$ polynomials.
@@
Figure 2 panel (b) shows that the largest errors occur almost periodically at a few places across $\Lambda=[0,1]$, confirming the equioscillation theorem.
@@ 6.2 Portfolio Allocation
We next consider a portfolio allocation problem calibrated to real data where $\bm{\theta}\in\mathbb{R}^{d}$ represents the weights on $d=10$ different asset classes. Namely, let $\mu\in\mathbb{R}^{d}$ and $\Sigma\in\mathbb{R}^{d\times d}$ be the mean and covariance matrix of the returns of the different asset classes. We fit these parameters to the monthly return data from Aug. 2014 to July 2024 using the Fama-French 10 Industry index dataset. [^2] https://mba.tuck.dartmouth.edu/pages/faculty/ken.french/data_library.html
@@
$\min_{\bm{\theta}}$ $\quad-\lambda_{1}\mu^{\top}\bm{\theta}+\lambda_{2}\bm{\theta}^{\top}\Sigma\bm{\theta}+\bar{\ell}_{1}(\bm{\theta}).$
@@
Ground truth $\bm{\theta}^{*}(\lambda)$ is computed over a $100\times 100$ grid.
We focus on bivariate-Legendre polynomials for our basis, scaled and shifted to be orthogonal to uniform distribution on $[0,1]\times[.2,1]$. We initialize with ALSP with 2 polynomials in each dimension, and iterate by adding one polynomial in each dimension, so that $q=4,9,16,25$, after which we stop.
@@
Figure 3: LSP and ALSP for Portfolio Allocation. Compares methods using the first $q=4,9,16$ bivariate-Legendre polynomials.
As in our previous experiment, we see that LSP rapidly converges to an irreducible error and then plateaus. By contrast, ALSP seems to make continued progress as we add polynomials. Both methods substantively outperform uniform discretization.
The improved performance over uniform discretization is partially attributable to the increased dimension of $\Lambda$ (because discretization suffers from the curse of dimensionality), but is also because (traditionally) discretization interpolates solutions in a piecewise constant fashion. In this example, $\bm{\theta}^{*}(\lambda)$ is very smooth. See Figure 6 in Appendix. Hence, polynomial can learn to interpolate values very fast, while uniform discretization needs a great deal more resolution.
6.3 Portfolio Allocation with Moderate Dimensional $\Lambda$
We next modify the portfolio allocation problem in Section 6.2 to include transaction costs relative to the current portfolio as follows:
$h(\theta,\lambda)\ =\ -\lambda_{1}\cdot\mu^{\top}\theta+\lambda_{2}\cdot\theta^{\top}\Sigma\theta+\|\theta-\lambda_{3:12}\|_{2}^{2}.$
Here $\lambda_{3:12}\in\mathbb{R}^{10}$ represents the current portfolio holding, and $\lambda\in\mathbb{R}^{12}$. We use the following basis
$\lambda_{(j\text{ mod }12)}\cdot\lambda_{2}^{\lfloor j/12\rfloor},\ \ j=1,\ldots,q,$
for each component of $\bm{\hat{\theta}}_{i}(\cdot)$, for $i=1,\ldots,10$, and let $q=12,24,36$ in our experiments. As $q\rightarrow\infty$, this basis contains the optimal solution path in its span for any choice of $\mu,\Sigma$ (c.f. Section B.2).
For simplicity, we ran LSP using torch.optim.LBFGS to solve an ERM version of the problem in Step 6 of Algorithm 1 on a training set of size $1000$, and present results for an independent validation set of size $1000$.
Discretization on a $12$-dimensional grid is computationally challenging. Hence, to assess our method we instead compute an ERM approximation to $\mathbb{E}\left[h(\bm{\theta}^{*}(\tilde{\lambda}),\tilde{\lambda})\right]$ over the training/validation sets, i.e., compute $\frac{1}{1000}\sum_{i=1}^{1000}h(\bm{\theta}^{*}(\lambda_{i}),\lambda_{i})$ by solving $1000$ separate optimization problems. This value is the dotted grey line in Figure 4. The smaller the gap, the closer $\bm{\theta}^{*}(\lambda_{i})$ is to the span of the basis.
Similar to previous experiments, as the basis size increases, performance improves, and even with a fairly small basis, the error is quite small.
Figure 4: Portfolio Allocation, $\dim(\Lambda)=10$. Comparing the ERM objective for varying $q$. Panel a) training, Panel b) validation.
@@ 7 CONCLUSION
We propose a new method for learning the optimal solution path of a family of problems by reframing it as a single stochastic optimization problem over a linear combination of pre-specified basis functions. Compared to discretization schemes, our method offers flexibility and scalability by taking a global perspective on the solution path. We prove that our problem satisfies a certain relaxed weak-growth condition that allows us to solve the single optimization problem very efficiently when using sufficiently rich bases. Theoretical results in special cases and numerical experiments in more general settings support these findings. Future research might more carefully examine the interplay between the parameterization of the family of problems and the choice of basis. One might also consider other universal function approximators (e.g. deep neural networks, forests of trees) within this context.
Acknowledgements
PG acknowledges the support of the NSF AI Institute for Advances in Optimization, Award 2112533. The authors thank the anonymous reviewers for their thoughtful and constructive comments.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Not Applicable]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Not Applicable]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ A.1 Proofs from Section 2
Suppose by contradiction that $c=0$. Then there exists a corresponding eigenvector $\bm{v}\in\mathbb{R}^{p}$ with $\bm{v}\neq\bm{0}$ such that
$\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\bm{\Phi}(\tilde{\lambda})^{\top}\bm{\Phi}(\tilde{\lambda})\right]\bm{v}=\bm{0}\implies\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\|\bm{\Phi}(\tilde{\lambda})\bm{v}\|^{2}_{2}\right]=0.$
This further implies that $\bm{\Phi}(\tilde{\lambda})\bm{v}=0$ almost everywhere. By the linear independence assumption in Assumption 2.2, we then must have $\bm{v}=\bm{0}$, but this is a contradiction since $\bm{v}$ is an eigenvector.
@@
$c\leq\sigma_{\max}\left\{\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\bm{\Phi}(\tilde{\lambda})^{\top}\bm{\Phi}(\tilde{\lambda})\right]\right\}\leq\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\sigma_{\max}\left\{\bm{\Phi}(\tilde{\lambda})^{\top}\bm{\Phi}(\tilde{\lambda})\right\}\right]\leq C,$
@@
$f(\bm{\beta},\lambda)~{}:=~{}h(\bm{\Phi}(\lambda)\bm{\beta},\lambda),\ \ \ \ F(\bm{\beta})~{}:=~{}\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[f(\bm{\beta},\tilde{\lambda})\right]=\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta},\tilde{\lambda})\right].$
Before proceeding, we establish the smoothness and strong convexity of these functions.
@@
Fix an arbitrary $\lambda\in\Lambda$. Then, for any $\bm{\beta}$, $\bar{\bm{\beta}}\in\mathbb{R}^{p}$, smoothness of $f(\cdot,\lambda)$ is certified by
$\|\nabla f(\bm{\beta},\lambda)-\nabla f(\bar{\bm{\beta}},\lambda)\|$ $=\|\nabla_{\bm{\beta}}h(\bm{\Phi}(\lambda)\bm{\beta},\lambda)-\nabla_{\bm{\beta}}h(\bm{\Phi}(\lambda)\bar{\bm{\beta}},\lambda)\|$
@@
$\ \leq CL\|\bm{\beta}-\bar{\bm{\beta}}\|$
for any $\bm{\beta}$, $\bar{\bm{\beta}}\in\mathbb{R}^{d}$. To obtain the same conclusions for $F(\cdot)$, we may take expectation over $\lambda\sim\mathbb{P}_{\lambda}$ on the above inequalities, and invoke the linearity of expectations as well as the property that
@@
Next, we verify the strong convexity of $F(\cdot)$. Fixing $\bm{\beta}$, $\bar{\bm{\beta}}\in\mathbb{R}^{d}$, for any $\lambda\in\Lambda$, define $\bm{\theta}(\lambda):=\bm{\Phi}(\lambda)\bm{\beta}$, $\bar{\bm{\theta}}(\lambda):=\bm{\Phi}(\lambda)\bar{\bm{\beta}}$. Then, by strong convexity of $h$, we have
$h(\bar{\bm{\theta}}(\lambda),\lambda)\geq h(\bm{\theta}(\lambda),\lambda)+\nabla_{\bm{\theta}}h(\bm{\theta}(\lambda),\lambda)^{\top}\left(\bar{\bm{\theta}}(\lambda)-\bm{\theta}(\lambda)\right)+\frac{\mu}{2}\|\bar{\bm{\theta}}(\lambda)-\bm{\theta}(\lambda)\|^{2}.$
@@
$F(\bar{\bm{\beta}})$ $=\mathbb{E}\left[h(\bm{\Phi}(\tilde{\lambda})\bar{\bm{\beta}},\tilde{\lambda})\right]$
$\ \geq\mathbb{E}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta},\tilde{\lambda})\right]+\mathbb{E}\left[\nabla_{\bm{\theta}}h(\bm{\theta}(\tilde{\lambda}),\tilde{\lambda})^{\top}\left(\bm{\Phi}(\tilde{\lambda})(\bar{\bm{\beta}}-\bm{\beta})\right)\right]+\frac{\mu}{2}\mathbb{E}\left[\|\bm{\Phi}(\tilde{\lambda})(\bar{\bm{\beta}}-\bm{\beta})\|^{2}\right]$
$\ =F(\bm{\beta})+\mathbb{E}\left[\bm{\Phi}(\tilde{\lambda})^{\top}\nabla_{\bm{\theta}}h(\bm{\theta}(\tilde{\lambda}),\tilde{\lambda})\right]^{\top}(\bar{\bm{\beta}}-\bm{\beta})+\frac{\mu}{2}(\bar{\bm{\beta}}-\bm{\beta})^{\top}\mathbb{E}\left[\bm{\Phi}(\tilde{\lambda})^{\top}\bm{\Phi}(\tilde{\lambda})\right](\bar{\bm{\beta}}-\bm{\beta})$
$\ \geq F(\bm{\beta})+\nabla F(\bm{\beta})^{\top}(\bar{\bm{\beta}}-\bm{\beta})+\frac{c\mu}{2}(\bar{\bm{\beta}}-\bm{\beta})^{\top}(\bar{\bm{\beta}}-\bm{\beta})$
@@
First, we prove that $\epsilon_{\rm sp}^{\ast}$ is attained at some $\bm{\beta}_{\rm sp}^{\ast}\in\mathbb{R}^{p}$. Observe that $\epsilon_{\rm sp}(\cdot)$ is continuous. Thus, its level set $M(1):=\{\bm{\beta}\in\mathbb{R}^{p}:\epsilon_{\rm sp}(\bm{\beta})\leq 1\}$ is closed. Furthermore, $M(1)$ is bounded. To see this, for any $\bm{\beta}\in M(1)$, by strong convexity of $F$,
$\frac{c\mu}{2}\|\bm{\beta}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}$ $~{}\leq~{}F(\bm{\beta})-F(\bm{\beta}_{\mathrm{avg}}^{\ast})$
$~{}=~{}\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta},\tilde{\lambda})-h(\bm{\Phi}(\tilde{\lambda})\bm{\beta}_{\mathrm{avg}}^{\ast},\tilde{\lambda})\right]$
$~{}\leq~{}\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta},\tilde{\lambda})-h(\bm{\theta}^{*}(\tilde{\lambda}),\tilde{\lambda})\right]$
$~{}\leq~{}\epsilon_{\rm sp}(\bm{\beta})$ (4)
@@
Since $\bm{\beta}$ is arbitrary, for any $\bm{\beta}$, $\bar{\bm{\beta}}\in M(1)$,
$\|\bm{\beta}-\bar{\bm{\beta}}\|~{}\leq~{}\|\bm{\beta}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|+\|\bar{\bm{\beta}}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|~{}\leq~{}\frac{2\sqrt{2}}{\sqrt{c\mu}}.$
Applying the Weierstrass Theorem, there exists $\bm{\beta}_{\rm sp}^{\ast}\in M(1)\subset\mathbb{R}^{p}$ s.t.
$\epsilon_{\rm sp}(\bm{\beta}_{\rm sp}^{\ast})~{}=~{}\epsilon_{\rm sp}^{\ast}.$
@@
$h(\bm{\Phi}(\lambda)\bm{\beta},\lambda)-h(\bm{\theta}^{*}(\lambda),\lambda)\ \leq\ \frac{L}{2}\|\bm{\Phi}(\lambda)\bm{\beta}-\bm{\theta}^{*}(\lambda)\|^{2}$ (5)